### v1

# Practice Makes Perfect: Planning to Learn Skill Parameter Policies

Abstract
One promising approach towards effective robot decision making in complex, long-horizon tasks is to sequence together parameterized skills. We consider a setting where a robot is initially equipped with (1) a library of parameterized skills, (2) an AI planner for sequencing together the skills given a goal, and (3) a very general prior distribution for selecting skill parameters. Once deployed, the robot should rapidly and autonomously learn to improve its performance by specializing its skill parameter selection policy to the particular objects, goals, and constraints in its environment. In this work, we focus on the active learning problem of choosing which skills to practice to maximize expected future task success. We propose that the robot should estimate the competence of each skill, extrapolate the competence (asking: “how much would the competence improve through practice?”), and situate the skill in the task distribution through competence-aware planning. This approach is implemented within a fully autonomous system where the robot repeatedly plans, practices, and learns without any environment resets. Through experiments in simulation, we find that our approach learns effective parameter policies more sample-efficiently than several baselines. Experiments in the real-world demonstrate our approach’s ability to handle noise from perception and control and improve the robot’s ability to solve two long-horizon mobile-manipulation tasks after a few hours of autonomous practice. Project website: https://planning-to-practice.github.io $*{}^{*}$qual contribution. Work done during internship at The AI Institute. Correspondence to {njk, tslvr}@csail.mit.edu.
I Introduction
Given the recent progress in robot skill learning and design [11, 59, 65, 39, 22], we are quickly approaching a future where robots will arrive at their deployment sites equipped with a library of general-purpose skills. Each robot will sequentially compose these skills in different ways to accomplish long-horizon tasks that will vary considerably between deployment sites. As the robot gathers experience during deployment, it should get better over time. In particular, the robot should learn to rapidly specialize its skills to the unique objects, goals, and constraints that it repeatedly encounters during deployment.
In this work, we consider skills that are continuously parameterized and we focus on parameter policy learning [40, 1, 15, 44, 20] as a mechanism for rapidly specializing skills. For example, a “pick” skill may be parameterized by a relative grasp and a “sweep” skill by a sweeping velocity (Figure LABEL:fig:teaser). Starting from general-purpose priors [54, 47, 7, 26], we want the robot to quickly learn specialized policies for selecting grasps, push velocities, and other skill parameters. Following previous work [33, 1, 52], we consider parameterized skills that are (extended) options [57]; each skill has an initiation condition, a parameterized controller, a termination condition, and a success condition. For example, a “place” skill can be initiated when the robot is holding an object and facing a surface; the skill terminates after the robot opens its gripper; and the skill is successful if the object is subsequently stably resting on the surface. Options are closely related to AI planning operators [33, 52] and we can leverage this relationship to efficiently plan a sequence of skills to reach a goal [28].
We consider parameter policy learning in the context of reset-free online learning [58, 38, 25, 42] where the robot alternates between solving given tasks (task time) and taking actions of its choosing (free time). For example, a given task might be to “clear objects off the table” (Figure LABEL:fig:teaser). We focus on free time and ask: how should the robot select actions so that, after learning parameter policies from the collected experience, the likelihood of solving given tasks in the future is maximized? This is an embodied active learning problem [16, 45, 36, 42], which is distinct from standard active learning [49] in that the robot must reason sequentially. For example, to collect one “sweep” data point in our experiments, the robot needs to execute up to 19 skills in sequence to reach a state where sweeping is possible (Figure LABEL:fig:teaser). This setting is also related to exploration in the reinforcement learning literature [29, 46, 13, 8, 2]; we consider baselines from that literature in experiments. Compared to end-to-end RL though, our setting has significantly more structure, which we can leverage to achieve much more sample-efficient learning, especially over long-horizon tasks.
To learn parameter policies through embodied active learning, we consider planning to practice parameterized skills. During free time, the robot repeatedly selects a skill, plans to a state where that skill can be initiated, practices the skill (selects continuous parameters to try), records the success condition outcome, and then updates its parameter policy accordingly. The central question is: how should the robot decide what skills to practice? One approach would be to practice the skill with the lowest competence [56, 14], that is, the skill that most often fails to achieve its success condition. But that skill may be impossible to improve or irrelevant to the given tasks.
We propose that the robot should instead practice the skill whose predicted competence improvement would maximally benefit the overall task distribution. Implementing this skill selection strategy requires three steps: estimating current skill competence; extrapolating the competence by predicting how much it would hypothetically improve through practice; and situating the competence in the task distribution by predicting how overall task success rates would hypothetically improve. We propose a Beta-Bernoulli time series model to estimate and extrapolate skill competence and use cost-aware AI planning [28] to situate the competence in the task distribution.
In experiments, we evaluate the extent to which our Estimate, Extrapolate & Situate (EES) approach enables the robot to make efficient use of its free time as measured by its success rate during task time. In three simulated environments, we compare to seven baselines and find that EES is consistently the most sample-efficient. We also implement EES in two real mobile manipulation environments using a Boston Dynamics Spot robot with an arm (Figures LABEL:fig:teaser, 1). In these environments, the robot plans and practices autonomously for several hours, coping with noise inherent to real-world perception and control, and rapidly improves its ability to solve long-horizon mobile-manipulation tasks.
II Problem Setting
This paper proposes a method for active practicing in the context of a robot system that has mechanisms for planning and learning. In this section, we describe our problem setting, including assumptions about the robot’s environment as well as minimal specifications for its planning and learning modules.
II-A Modelling the World
We assume that the robot and its environment are modelled as a goal-based Markov Decision Process with object-oriented states [17] and parameterized actions [40, 1]. States are factored into objects and their continuous features. For example, consider the Ball-Ring environment shown in Figure 1. The ball, ring, table, floor, and robot itself are objects, and their features include, for example, gripper joint value (for the robot) and xyz position (for other objects). Of course, a real-world robot cannot perceive such features directly, so we assume that the robot is equipped with a perception system that can construct a fully-observed state $x_{t}\in\mathcal{X}$ from raw sensory observations at each time step $t\in\mathbb{Z}^{+}$. This model does not account for the perception noise that exists in the real world, but in experiments, we find that our approach is reasonably robust to that noise (Section IV).
The action space of the MDP is defined by a set of parameterized skills that have continuous parameters. It is often convenient to define object parameters as well, but for the purpose of this work, these play a minor role, so we treat them as part of the skill except where otherwise noted. For example, in the Ball-Ring environment, Place($\circ$) is one skill $u\in\mathcal{U}$, where $\circ$ denotes a placeholder for a continuous parameter, and Place($\circ$) is another $u^{\prime}\in\mathcal{U}$. The continuous parameters for both skills are xy relative offsets between the gripper and target surface (the height and gripper orientation are fixed).
To define skills formally, we use an extension of the options framework [57]. A parameterized skill $u\in\mathcal{U}$ is given by a tuple $(I,\Theta,\mu,\beta,J)$ where $I:\mathcal{X}\to\{0,1\}$ characterizes states where the skill can be initiated, $\Theta\subseteq\mathbb{R}^{m}$ is the set of possible continuous parameters, $\mu(x,\theta)$ is a low-level controller that takes a state $x$ and continuous parameters $\theta\in\Theta$ as input, $\beta:\mathcal{X}\to\{0,1\}$ is a termination condition, and $J:\mathcal{X}\to\{0,1\}$ is a success condition indicating whether the skill has achieved its intended outcome in the terminal state. A skill with parameters assigned is treated as an atomic action. After an action $a_{t}\in\mathcal{A}$ is executed, the environment advances according to an unknown transition distribution $x_{t+1}\sim P(\cdot\mid x_{t},a_{t})$.
The robot is tasked with achieving particular goals. Each goal is sampled from a task distribution $g\sim P(\cdot\mid x_{0})$. For example, in the Ball-Ring environment, the goal might be that the ball is stably at rest on the table. We do not assume direct access to the goal distribution; instead, the robot receives goals from a human gradually during learning II-C. Formally, a goal is a binary classifier over states $g:\mathcal{X}\mapsto\{0,1\}$ where 1 indicates that a state is within the goal set. We refer to a combination of an initial state $x_{0}$ and goal $g$ as a task. Solving a task entails taking actions to reach a state $x_{t}$ where $g(x_{t})=1$ from $x_{0}$ within a maximum time-step horizon $H_{\text{eval}}$.
II-B Planning to Solve Tasks
Given a task, the robot will plan to generate actions that are likely to accomplish the goal from the initial state. Following previous work [55, 1, 51, 35], we decompose planning into two levels: skill sequencing and continuous parameter selection. Skill sequencing consists of generating a skeleton, e.g., ((((MoveTo($\circ$), Pick(ball, $\circ$), MoveTo($\circ$), Place(ball, $\circ$))normal-))). Given a skeleton, we select continuous parameters using parameter policies $\theta\sim\pi_{u}(\cdot\mid x)$. Since parameter selection is conditioned on the state $x$, and since we are not assuming a known transition distribution [51], we sample and execute each skill greedily. If the skill terminates and does not meet its success condition, we replan. See Algorithm 1 for a summary.
Figure 1: Running example: Ball-Ring environment. The goal is to put the ball on the table. The robot should learn that (1) the ball cannot be placed directly because it will roll off the slanted table; (2) the ring can only be placed on the left side because the right side is smooth (shown in the top-right corner); (3) placing the ring on the table and then placing the ball inside the ring is the best way to accomplish the goal.
The robot will learn parameter policies through online experience (Section  II-C). We assume that each skill $u$ is accompanied by a parameter prior $\pi^{0}_{u}$ to be used before any parameter policies have been learned. For example, a pick skill may have an associated grasp sampler that provides valid grasps some percentage of the time. At first, the robot uses the parameter priors to select parameters ($\pi_{u}=\pi^{0}_{u}$), but as the robot collects online experience, it will learn to improve the parameter policies with respect to the environment and task distribution. For example, the robot should learn grasp samplers that are specialized to the objects in the environment that need to be manipulated.
How can we generate skeletons to maximize the probability that sampling will succeed? Towards answering this question, we introduce the notion of skill competence.
Definition 1 (Skill Competence).
The competence $c_{u,\pi}$ of a skill $u$ with current parameter policy $\pi_{u}$ is the expected success $\mathop{\mathbb{E}}[J_{u}(X_{t+1})\mid I_{u}(X_{t})]$, where $X_{t}$ is a random variable for the state before skill execution, $A_{t}$ is a r.v. for the action generated from $\theta\sim\pi_{u}(\cdot\mid X_{t})$, and $X_{t+1}\sim P(\cdot\mid X_{t},A_{t})$.
For example, if Pick(ball, floor, $\circ$) successfully grasps the ball from the floor 80% of the time, the competence would be 0.8. Note that competence is defined in terms of the current parameter policy, and that the distribution of $X_{t}$ is induced by the overall planning procedure and the task distribution. In practice, skill competences are unknown and must be estimated from data (Section III).
To establish a relationship between skill competence and full skeleton success, we introduce a strong assumption:
Assumption 1.
Success $J_{u}(X_{t+1})$ is independent from the state $X_{t}$ conditioned on the initiation condition $I_{u}(X_{t})=1$.
In other words, the success rate of a skill is the same for all states in its initiation set. This assumption has been previously considered in different forms [33, 1], but it does not always hold in practice. For example, the specific grasp of an object may influence the success rate of placing. We can mitigate this by replanning, but to fully remove the assumption, we would need task and motion planning [55, 24] or automated skill partitioning [33, 1], which we leave to future work.
We can now revisit the problem of generating a skeleton that has the maximum likelihood of success. Given Assumption 1, we want to find a skeleton $(u_{0},\dots,u_{n})$ with three properties: (1) $\prod_{i=0}^{n}c_{i}$ is maximal, where $c_{0},\dots,c_{n}$ are the corresponding competences for the skills in the skeleton; (2) the initiation and success conditions for subsequent skills chain together (see [33] for a formal definition); and (3) the goal is achieved. Following previous work [1, 33, 51, 52], we take advantage of the close relationship between (parameterized) options and AI planning operators to generate skeletons that satisfy conditions (2) and (3). Previous work has considered how to learn these operators automatically; we manually specify them for this work. To satisfy condition (1), we associate a cost of $-\log(c)$ to the respective operator and use an off-the-shelf AI planner [28] to find a minimal cost (maximum likelihood) skeleton. See Appendix -A for further details.
654321 Input: Current state $x$ and goal $g$. Generate a skeleton $(u_{0},\dots,u_{n})$ to $g$ from $x$. For $i=0,\dots,n$:  Sample $\theta\sim\pi_{u_{i}}$ and execute $u_{i}(\theta)$.  Perceive and update the current state $x$.  If $J_{u_{i}}(x)\neq 1$, repeat from line 2 (replan).
Algorithm 1 Planning and Execution
7654321 Initialize parameter policies $\Pi=\{\pi^{0}_{u}:u\in\mathcal{U}\}$. Repeat:  If a human has given a goal $g$:   Plan and execute to $g$ with Algorithm 1.  Else:   Select and execute actions of the robot’s choice.  Update $\Pi$ every $m$ iterations.
Algorithm 2 Online Learning Paradigm
II-C Online Learning Paradigm
We want the robot to get better at solving tasks over time. We consider a reset-free online learning [58, 38, 25] paradigm where the robot is sometimes given a task to solve and otherwise given free time during which it should autonomously learn to improve. The key question is: what should the robot do during free time to get better at solving tasks?
We assume that the skills themselves are fixed (e.g., for safety reasons), but the parameter policies can change. The robot should therefore use its free time to improve its parameter policies, specializing the given parameter priors to the particular objects, goals, and constraints in its environment. This setup is summarized in Algorithm 2. Note that this setup is fully autonomous; the environment is not reset. Our main interest is Line 6: how should the robot choose actions to gather data for improving its parameter policies?
III Planning to Learn
7654321 Input: Current parameter policies $\Pi$. Select a skill $u\in\mathcal{U}$ to practice (see Algorithm 4). Plan to $I_{u}$ using Algorithm 1 with $\Pi$. Practice the skill $u$ one time:  Sample parameters $\theta$ from an explore policy $\pi^{+}_{u}$.  Execute $u(\theta)$ and record the transition. Repeat from line 2 until free time expires.
Algorithm 3 Planning to Practice
We propose that the robot should spend its free time planning to practice skills. In particular, we commit to the meta-strategy shown in Algorithm 3, where the robot repeatedly selects a skill to practice, plans to satisfy that skill’s initiation condition, samples parameters from an explore parameter policy, executes the action, and records the result. In using this meta-strategy, we make two assumptions.
Assumption 2.
For $x\in\mathcal{X}$ and $u\in\mathcal{U}$, there exists a sequence of actions that reach $I_{u}$ from $x$ with nonzero probability.
In other words, it is not possible to get permanently “stuck” during online learning. This assumption can be weakened if certain skills do not need to be practiced infinitely often.
Assumption 3.
(Informal) Parameter priors have support over good parameter choices.
For efficiency (and perhaps safety) purposes, we will not permit the robot to sample arbitrarily from skill parameter spaces; we therefore assume that the priors are sufficiently broad to enable learning. Given these assumptions, we are left with three decisions:
  1. 1.
    How should we decide what skills to practice?
  2. 2.
    What explore parameter policies $\pi^{+}_{u}$ should we use?
  3. 3.
    How should we update the parameter policies?
In this work, we choose to focus on the first question and draw on existing techniques to answer the second two. See Figure 2 for an overview of the full pipeline.
III-A Selecting Skills to Practice
Given the relationship between competence and task success, a natural answer to the first question would be to practice the skill with the lowest current competence. However, there there are two major issues with this “Fail Focus” strategy. First, a low-competence skill may be impossible to improve. For example, in the Ball-Ring environment, the Place(ball, table, $\circ$) skill is bound to fail since the table is slanted (as seen in Figure 1). Second, even if a low-competence skill could be improved, the skill may be less critical for the task distribution than others. In the worst case, Fail Focus may cause the robot to spend all its free time attempting to improve an impossible skill that is of no consequence to any given task.
A better skill selection strategy would be more directly tied to our real objective: to efficiently and effectively solve the tasks given to the robot. We consider a close proxy to this real objective. Given parameter policies $\Pi=\{\pi_{u}:u\in\mathcal{U}\}$ and a task $(x_{0},g)$, let $J_{\text{task}}(\Pi,x_{0},g)$ be the probability that planning succeeds without replanning. From Section II-B, we have that $J_{\text{task}}(\Pi,x_{0},g)=\prod_{i=0}^{n}c_{i}$, where $c_{i}$ is the competence of the $i^{\text{th}}$ skill in the skeleton generated for the task. Given a task distribution $(X_{0},G)$, our overall objective is to learn parameter policies that maximize:
$J_{\text{tasks}}(\Pi)\triangleq\mathop{\mathbb{E}}_{X_{0},G}[J_{\text{task}}(\Pi,X_{0},G)],$ (1)
In other words, the expected probability that planning succeeds without replanning over the task distribution.
We propose to practice the skill whose predicted improvement would maximally increase $J_{\text{tasks}}(\Pi)$. In other words, we will practice the skill with the greatest expected improvement to the overall distribution of human-given tasks. We do this in three key steps (Algorithm 4). For each skill, we: (1) Estimate the competence: compute the current competence of the skill from data; (2) Extrapolate the competence: predict how the competence of the skill would change if it were practiced once more and then its parameter policy was updated; (3) Situate the competence: predict how the overall success rate on the task distribution would change given the extrapolated skill competence. This approach resolves the issues with Fail Focus: by extrapolating, we avoid practicing impossible or plateaued skills; and by situating, we avoid practicing irrelevant or unimportant skills. We now describe these steps in detail.
321 For each $u\in\mathcal{U}$ with current parameter policy $\pi_{u}$:  Estimate the current competence $c_{u,\pi}$.  Extrapolate: predict $c_{u,\pi^{\prime}}$, the competence after
4  practicing $u$ and updating $\pi_{u}$ to $\pi_{u}^{\prime}$.  Situate the competence in the task distribution,
5  computing $J_{\text{skill}}(u)\triangleq J_{\text{tasks}}(\Pi-\{\pi\}\cup\{\pi^{\prime}\})$. Return $\operatorname*{argmax}_{u}J_{\text{skill}}(u)$ for practice.
Algorithm 4 Selecting a Skill to Practice
Figure 2: Pipeline overview. (1) During free time, the robot repeatedly selects skills to practice. Here, Place($\circ$) is selected because it maximizes $J_{\text{skill}}$ (Algorithm 4). (2) The robot plans to satisfy the initiation condition of the skill and then selects a continuous parameter to practice (Algorithm 3). (3) The resulting success or failure of the skill is used to improve the parameter policy (Section III-C).
III-A1 Estimating Skill Competence
Our first task is to estimate the current competence of a skill based on the transitions that have been collected thus far. To estimate competence, we propose a graphical model that explicates the relationship between competence, transitions, and learning.
Recall that parameter policies are updated periodically (Algorithm 2); we refer to each period with the same parameter policy (i.e., before an update is made) as a learning cycle. Thus, the robot’s free time is composed of a series of learning cycles. Let $S_{i,k}$ be a binary random variable for the result of the success condition on the $k^{th}$ usage of the skill in learning cycle $i$ [^1] We have not yet defined the explore policy, but it is important here that only “exploit” samples are used to estimate competence; see Section  III-B. Note that different skills are selected for practice each cycle, so the number of skill usages (maximum $k$ values) varies and can be zero. In practice, this number is low, so we cannot reliably estimate the competence of a skill based on data from the current cycle alone.
Let $C_{i}$ be a random variable for the skill’s competence during learning cycle $i$ (before it has re-learned using the data from that cycle). To estimate skill competence, we wish to know $P(C_{k}\mid S_{1,1},S_{1,2},\dots,S_{1,n_{1}},S_{2,1},\dots,S_{k,n_{k}})$, that is, the conditional distribution of competence now given all observations up until now. We consider a Bayesian time series model with a joint distribution that factorizes as follows:
$P(S_{1:n_{1}},...,S_{k:n_{k}},...,C_{1:N_{\text{cycle}}})=\prod_{i}P_{i}(C_{i})\prod_{k}P(S_{i,k}\mid C_{i})$
where $P(S_{i,k}\mid C_{i})$ is the observation model and $P_{i}(C_{i})$ is a cycle prior. For the observation model, we use a Bernoulli:
$P(S_{i,k}=1\mid C_{i}=c)=c.$
For the cycle priors, we use Beta distributions, since the Beta is the conjugate prior of the Bernoulli. Let us first consider $P_{0}(C_{0})$, the prior competence before any observations have been made. We define one prior for all skills that has high mean but large variance (in experiments, $\text{Beta}(10,1)$), reflecting our weakly held expectation that parameter priors $\pi^{0}_{u}$ will be generally good. This also introduces a form of optimism, which can be helpful for exploration [4, 41].
For subsequent cycle priors $P_{i}(C_{i})$, we assume that a skill’s competence is some function of the size of the dataset used to learn that skill’s parameter policy. Note that we are not positing a general relationship between data count and competence. One skill may always have perfect competence; another may always have zero competence; and a third may improve as data increases. Let $f_{\phi}:\mathbb{Z}_{\geq 0}\to\text{Beta}(\alpha,\beta)$ be a competence model where the input is the number of data used for learning, the output is a Beta distribution over competence and $\phi$ indicates that the function belongs to a hypothesis class $\mathcal{F}$. The cycle prior $P_{i}(C_{i})$ is given by $f_{\phi}(m_{i})$ where $m_{i}$ is the number of data collected for the skill through cycle $i$. For example, in Ball-Ring, a good competence model for the Place(ball, $\circ$) skill would output a near-zero Beta for any input, because no amount of data can improve the skill. On the other hand, a competence model for Place(ring, $\circ$) should output Beta distributions with increasing modes, since that skill can improve with practice.
To estimate the current skill competence, we need to infer $C_{i}$ for all $i$ and fit $\phi$ for $f_{\phi}$. We considered two approaches: a principled expectation-maximization (EM) approach, and a much simpler sliding-window-based approach. In preliminary experiments, we found the simpler approach to perform at least as well as EM, and its behavior was much easier to interpret, so we used it for our main experiments. See Appendix -B.
III-A2 Extrapolating Skill Competence
Given the competence model $f_{\phi}$ fit during estimation, extrapolation is straightforward: we can simply evaluate $f_{\phi}(m+1)$ to predict how the skill competence would change if we collected one more data point of practice, where $m$ is the number of data seen so far. Let $\hat{c}$ denote the mode of $f_{\phi}(m+1)$, i.e., the most likely next competence. Here we assume that a skill’s competence never gets worse with learning: $\forall\phi$, if $m^{\prime}>m$, then $\mathbb{E}[f_{\phi}(m^{\prime})]\geq\mathbb{E}[f_{\phi}(m)]$. We can enforce this assumption by choosing $\mathcal{F}$ appropriately. This assumption may not always hold in practice, but for the purpose of extrapolation, it is important that the agent be optimistic and not deliberately avoid collecting additional data for a skill.
III-A3 Situating Skill Competence
Our final step is to predict the expected improvement to the overall task distribution given the extrapolated competence. Let $\Pi^{\prime}$ be the set of current parameter policies $\Pi$, but with the parameter policy $\pi_{u}$ for the current skill under consideration $u$ replaced with $\pi^{\prime}_{u}$, a hypothetical policy that would result from practicing $u$ once more and re-learning. We wish to compute $J_{\text{tasks}}(\Pi^{\prime})$ (Equation 1), the expected probability that planning would succeed (without replanning) over the task distribution.
To compute $J_{\text{tasks}}(\Pi^{\prime},x_{0},g)$ for a given task $x_{0},g$, we need not know $\pi^{\prime}_{u}$ itself, but only the competence of $\pi^{\prime}_{u}$, which we have computed by extrapolating. To complete $J_{\text{tasks}}(\Pi^{\prime})$, we need to take an expectation over the task distribution. As mentioned in Section II, we do not assume that the robot has direct access to the task distribution; instead, we collect the states and goals used to query the planner, including when replanning is triggered (Algorithm 1) and use that empirical task distribution to approximate $J_{\text{tasks}}(\Pi^{\prime})$.
III-B Explore Parameter Policies
After we have selected a skill to practice and planned to satisfy its initiation condition, we must decide what parameters to use (Algorithm 3). We can view this parameter selection problem as a contextual bandit with infinite arms [6, 37]. Here, the context is the current state $x_{t}$, the actions are parameters $\theta$, and the reward is 1 if the success condition passes and 0 otherwise. To balance exploration and exploitation, we use an epsilon-greedy policy:
$\pi^{+}_{u}=\epsilon\pi^{0}_{u}+(1-\epsilon)\pi_{u}.$
Other choices are possible; we use this simple approach to maintain focus on the skill selection problem.
III-C Learning to Improve Parameter Policies
To complete our approach for planning to practice parameterized skills, we must now determine how the collected experience can be used to improve the parameter policies. Recall that for each skill $u$, we are given a parameter prior $\pi^{0}_{u}$ and we wish to learn an improved parameter policy $\pi_{u}$. Recall also that we have recorded transition data, which we can partition by skill and label according to whether the $J$ check for that particular skill passed: $\mathcal{D}_{u}=\{((x_{t},a_{t}),J_{u}(x_{t+1})):a_{t}\text{ uses skill }u\}$. Given these data, many approaches are possible. Our approach is to learn an implicit (energy) function $E_{u}:\mathcal{X}\times\mathcal{A}\to\mathbb{R}_{\geq 0}$ to define the parameter policy:
$\pi_{u}(\theta\mid x)\propto\pi^{0}_{u}(\theta\mid x)E_{u}(x,u(\theta)).$
Specifically, we train small neural network classifiers to minimize binary cross entropy loss and then use the classifier log probabilities for $E_{u}$. In contrast to the explore policy, the parameter policy $\pi_{u}$ is meant to exploit, so we select parameters via $\operatorname*{argmax}_{\theta}{\pi_{u}(\theta\mid x})$. In practice, we sample 100 candidates from the prior $\pi^{0}_{u}(\theta\mid x)$ and select the maximum.
For real-robot experiments, it is essential that parameter policies can be learned from very little data. Beyond using good parameter priors, we take two additional steps for data efficiency. First, we share neural network weights between parameter policies that have the same “parent” skill but different object parameters (e.g., Place($\circ$) and Place($\circ$)). Second, we perform feature engineering for learning by mapping the full state and action to a low-dimensional vector that is input to $E_{u}$. After applying these features to the data in $\mathcal{D}_{u}$, we are left with a standard binary classification dataset. See Appendix -E for details of the feature mapping as well as the training and use of $E_{u}$.
IV Experiments
Our experiments are designed to empirically answer the following questions about our approach (EES):
  • Q1.
    To what extent does EES choose skills for practice that lead to improvements in task distribution success rate, especially compared to alternative approaches?
  • Q2.
    How sample efficient is EES compared to alternatives?
  • Q3.
    To what extent is EES aware of the task distribution?
Figure 3: Simulation results. Percentage of evaluation tasks solved vs. number of online transitions collected for all approaches in all simulated environments. Solid lines represent means and shading represents standard error across 10 seeds.
Environments. We now provide high-level environment descriptions with details in Appendix -F. We use three simulated environments of varying complexity and two real-robot analogs to simulated environments. For details on our real-robot setup, see Appendix -D. See also the supplementary material for time-lapse videos of the real robot practicing skills and learning over time.
  • $\bullet$
    Light Switch (Simulated): A toy 1D grid environment. The robot starts in the leftmost room (grid cell) and must switch on a light in the rightmost room. The light is controlled by a dial in the same room that must be precisely actuated. The robot has skills to move left or right, turn the dial to a sampled setting, and also try to “jump” from a particular room all the way to the final room with the dial, though this jump skill is impossible and always fails. We use a grid size of 25 rooms in our main experiments.
  • $\bullet$
    Ball-Ring (Simulated): A simulated version of the environment depicted in Figure 1. To add complexity, the simulated version features multiple tables, some of which are slanted and partially smooth (as shown in the figure) and others that are standard flat surfaces. The relative locations of the smooth patches vary between tables. The robot does not initially know that placing the ball on a slanted surface will fail, nor does it know that placing the ring on the smooth part of a slanted table will fail with high probability. The robot has skills for moving, picking, and placing the ball and ring.
  • $\bullet$
    Ball-Ring (Real): The real version of the previous environment. See Figure 1 and the supplementary video.
  • $\bullet$
    Cleanup Playroom (Simulated): A simulated version of the environment depicted in Figure LABEL:fig:teaser. The robot is tasked with cleaning up a child’s playroom by putting two toys into a bin. The toys start out atop a table that may be blocked by a chair. There is also a brush on the floor. The robot again has skills for moving, picking, and placing, but also for grasping and dragging the chair, grasping and dumping out the bin, and sweeping toys from the table into the bin. The success of sweeping depends on the relative positions of the toys and bin and the sweeping velocity.
  • $\bullet$
    Cleanup Playroom (Real): The real version of the simulated Cleanup Playroom environment. See Figure LABEL:fig:teaser and the supplementary video.
Approaches. We now briefly describe all of the approaches that we evaluate. See Appendix -E for details. The first five approaches are alternative instantiations of planning to practice (Algorithm 3); the last three do not use that meta-strategy.
  • $\bullet$
    Estimate, Extrapolate, Situate (EES): Our main approach.
  • $\bullet$
    Fail Focus: Practices the skill with the lowest current estimated competence.
  • $\bullet$
    Competence Gradient: Inspired by [56, 5, 13], practices the skill with the highest expected competence gain (i.e., difference between current and extrapolated competence).
  • $\bullet$
    Skill Diversity: Practices the least-practiced skill.
  • $\bullet$
    Task-Relevant: Practices a randomly selected skill among those that have been previously included in some maximum-likelihood task plan.
  • $\bullet$
    Task Repeat: Samples a task from the empirical task distribution and plans to the task goal. If the task goal is already reached, plans instead to the task initial state.
  • $\bullet$
    Random Skills: Repeatedly executes random skills selected from those whose initiation conditions are satisfied in the current state. Does not plan.
  • $\bullet$
    MAPLE-Q: Uses hierarchical reinforcement learning to learn to select both skills and parameters. Does not plan. This approach is a modified version of MAPLE [44], which was developed for a more difficult problem setting. Our modifications are meant to give the approach access to the same prior knowledge as the other approaches (e.g., explicit option definitions and parameter priors), but a direct comparison remains challenging. See Appendix -E for further discussion and details.
# Steps S1 S2 S3 S4 S5 Mean
0 0.0 0.0 0.0 0.0 0.0 0.0
120 0.0 0.33 0.0 0.67 0.0 0.2
240 1.0 0.67 1.0 0.67 0.67 0.8
TABLE I: Ball-Ring real-robot results. Steps are online transitions; S1 = Seed 1. Entries are fractions of 3 evaluation tasks solved.
Experimental Setup. For all simulated environments, we run 10 random seeds of each approach. Task generation, parameter prior sampling, tie-breaking during planning, and stochastic environment transitions all vary between seeds. Within each environment, the number of free periods, the number of steps within each free period, and the evaluation horizon $H_{\text{eval}}$ is the same for all approaches (see Appendix -F). After every free period, each approach is evaluated on 10 randomly sampled “evaluation tasks” with held-out initial states. All simulation experiments were conducted on a quad-core Intel Xeon Platinum 8260 processor with a 192GB RAM limit. For real-robot environments, we run our main approach only (due to the intense time and resource requirements) and use 5 random seeds. Each seed represents a fully independent learning trial. We load model checkpoints after learning and use 3 tasks to evaluate performance. For all experiments, our key quantitative measure is the robot’s task success rate under the evaluation horizon.
Results and Analysis. Figure 3 shows a plot of the evaluation success rate of all approaches in all simulated environments. EES is consistently the most sample efficient, achieving higher success rates after fewer online transitions than the baselines. Fail Focus falters not only because it focuses on impossible skills, like jumping in Light Switch or placing the ball on the table in Ball Ring, but also because it is not sufficiently situated in the task distribution. For example, there are no impossible skills in Cleanup Playroom, but EES still outperforms Fail Focus. Competence Gradient is competitive with Fail Focus but similarly lacks situatedness. The poor performance of Skill Diversity, Task-Relevant, Task Repeat, and Random Skills underscore the importance of directed active practice in these long-horizon environments with many possible skills. Like the Random Skills baseline, MAPLE-Q fails to solve any evaluation tasks, which is not surprising given the highly limited number of online transitions that we are considering; previous related work [44] learns from multiple orders of magnitude more data. We verified that our implementation does well in far simpler environments with far more data; see Appendix -E for further discussion.
# Steps S1 S2 S3 S4 S5 Mean
0 0.0 0.67 0.33 0.67 1.0 0.53
120 0.67 1.0 1.0 1.0 1.0 0.93
240 1.0 1.0 0.67 1.0 1.0 0.93
TABLE II: Cleanup Playroom real-robot results. See Table I caption for details.
Tables I and II show EES evaluation results for the real-robot environments. Our approach is able to improve its performance after 120 and 240 real-world skill-executions respectively. Each seed took 1-3 hours of real robot time; see the supplementary video. These results are especially noteworthy given the non-trivial noise in perception and control, leading to both false positives and false negatives in the skill datasets ($\mathcal{D}_{u}$) collected for parameter policy learning.
Towards answering Q3 (to what extent is EES aware of the task distribution?), we conducted an additional experiment in the simulated Cleanup Playroom environment. Rather than giving the robot the goal of putting both toys into the bin, we instead gave the goal of putting one toy in the bin. As shown in Table III, this task distribution change led to a corresponding change in practicing behavior: rather than practicing sweeping, the robot instead practiced dropping the toy into the bin directly. This finding matches intuition: sweeping is unnecessarily complicated when only one object needs to be stowed (a pick-and-place strategy is better), but worthwhile when two objects can be stowed with one sweep.
Goals Pick(brush) Drop(toy, bin) Sweep(...)
Both toys 217.1 0.0 110.2
One toy only 88.4 305.9 0.0
TABLE III: Skills practiced with varying task distributions. The entries are the total number of times EES chooses to practice a select number of skills in the simulated Cleanup Playroom environment (averaged over 10 seeds). See text for discussion.
V Related Work
V-A Exploration in Reinforcement Learning
The problem of sequentially selecting actions that lead to efficient learning is central to exploration in reinforcement learning [29, 46, 13, 8, 2]. One important difference between our setting and RL is that we do not have a temporal credit assignment problem: given skill success conditions and Assumption 1, each parameter policy learning problem is self-contained. Our skill selection problem is therefore related to exploration in multi-armed bandits [9], but different still, since selecting a skill to practice does not lead to a task reward, but rather, to a data point that can be used to improve the parameter policy for that skill. Note that the inner problem of choosing parameters for a selected skill is a bandit problem (with infinite arms) [10] as explained in Section III-B, but our primary interest is the outer skill selection problem.
Within the RL literature, the most related work to ours is that of Stout and Barto [56], Baranes and Oudeyer [5] and Colas et al. [13], who each consider a form of competence progress to guide exploration. Compared to our approach, these previous works estimate and extrapolate competence, but they do not situate the competence in a task distribution. Their motivation is different from ours—they assume that a task distribution is not known and consider the problem of deriving intrinsic motivation in the absence of goals [48]. The Competence Gradient baseline in our experiments is inspired by these works and confirms the importance of situating competence in our setting.
V-B Parameter Policy Learning in RL
The problem of learning skill parameter policies has also been considered in the RL literature, for example, in Parameterized Action MDPs (PAMDPs) [27, 40, 15, 44]. In addition to facing the challenge of temporal credit assignment, these works typically do not assume a given method for discrete skill sequencing and instead need to learn a high-level “manager” policy in addition to the “worker” parameter policies. Altogether, this represents a much harder problem setting than ours, and the sample complexity of current techniques remains prohibitively high for the kind of rapid skill specialization we consider here (for example, see our MAPLE-Q baseline [44]). Assuming that the environment can be automatically reset is the norm in this literature, with some notable exceptions [58, 38, 25]. Work by Ames et al. [1] is a step toward bridging the RL setting and our setting; they automatically derive AI planning operators from parameterized option specifications. Exploration is not a central consideration in these works (but see [14, 3]).
Recent work in RL also considers specializing (fine tuning) skills through online learning [26, 62], starting from generic (pre-trained) distributions. Other recent work by Fang et al. [19] is another example of planning to practice. Their planning uses a learned latent subgoal space, rather than options and AI planners. They learn goal-conditioned policies that are analogous to our parameter policies. Active learning at the skill level is not a primary focus; their approach is similar to our Task Repeat baseline (but with environment resets instead of reset-free learning). Follow-up work [20] considers active learning more centrally and proposes a method for generating subgoals for online learning using a diversity-based metric.
V-C Learning Samplers for Task and Motion Planning
In the context of the task and motion planning (TAMP) literature [24], our parameter policies can be seen as samplers for refining skeletons generated by task planning. TAMP approaches typically do not make Assumption 1 and instead sample parameters contingent on the entire skeleton (e.g., selecting grasp parameters that enable future constrained placements).
Several works have considered learning samplers for TAMP [30, 12]. Silver et al. [52] learn samplers from an offline demonstration dataset. The details of our neural-network learning over object-centric states are most similar to theirs. Other recent work has considered learning samplers with diffusion models [43, 64]. Most relevant of these is the work by Mendez-Mendez et al. [42], who consider diffusion-based sampler learning for TAMP in an embodied lifelong setting. However, in that work, the robot is not given free time; it remains in task time throughout online learning. Additionally, the agent does not have separate exploration and exploitation samplers, but rather only an exploitation sampler. Wang et al. [60] consider active sampler learning for TAMP with a focus on the inner bandit problem of selecting parameters to practice for a given skill. In principle, their parameter selection method could be swapped in for our epsilon-greedy approach.
In the TAMP literature, the work by Noseworthy et al. [45] is another instance of active learning. They learn to predict whether a skeleton is feasible [61, 18, 63], i.e., whether there exists continuous parameters that would achieve the goal (typically in a deterministic setting). Future work could combine active feasibility prediction with our active parameter policy learning as a path toward removing Assumption 1 and scaling to more geometrically complex environments.
VI Limitations and Future Work
In this work, we proposed Estimate, Extrapolate & Situate (EES) as a method for planning to practice parameterized skills. We found that simulated and real robots using EES are able to rapidly and continually improve their parameter policies with respect to human-given task distributions. Our real-robot results are particularly noteworthy as instances of reset-free online learning in challenging, long-horizon mobile manipulation environments.
There are several limitations of the present work and of EES as a general method. For the sake of rapidly learning on a real robot, we started with a considerable amount of prior knowledge: known object (feature) detectors, fully-specified parameterized skills (and operators for planning), low-dimensional feature selectors for parameter policy training, and good parameter priors. Previous work has considered learning each of these components (e.g.,  [52, 54]), but doing so may require significantly more data than what we considered here. We also made Assumptions 1-3, which are strong, and while we need not satisfy them completely to attain good performance, they remain worthy of further scrutiny. Our relatively naive treatment of noise and our assumption of full observability are also clearly limiting. Furthermore, our commitment to planning to practice (Algorithm 3) is perhaps overly myopic: better strategies might anticipate that practicing one skill enables quickly practicing another, reasoning over sequences of practice attempts. Finally, this work presupposes that robots should be practicing and learning during deployment. The extent to which this is true depends greatly on the nature of the deployment and the constraints under which the robot is allowed to practice.
One future direction that could address multiple limitations simultaneously would be to give the agent access to a simulator. Even if the simulator were a coarse approximation of the real world, the robot could nonetheless use it to bootstrap real-world practice time. The same simulator could be used for reasoning about potentially irreversible actions before executing them in the real world, and for integrated task and motion planning (TAMP) [24] towards removing Assumption 1. Leveraging TAMP would also be a step toward more principled planning in stochastic [50] and partially-observable environments [23].
Acknowledgments
We gratefully acknowledge support from NSF grant 2214177; from AFOSR grant FA9550-22-1-0249; from ONR MURI grant N00014-22-1-2740; from ARO grant W911NF-23-1-0034 and from the MIT Quest for Intelligence. Nishanth, Tom, and Willie are supported by NSF GRFP fellowships. We thank Will Shen for feedback and suggestions on an early paper draft, as well as invaluable help with creating the accompanying website. We also thank Russell Mendonca for helpful early discussions, especially with respect to the MAPLE-Q baseline. We are grateful for helpful discussion, brainstorming, and support from Stefanie Tellex, Ashay Athalye, Tushar Kusnur, Jiuguang Wang, Gustavo Goretkin, Andrew Messing, Joe St. Germain and others at the AI Institute. We thank Hannah Blumberg for invaluable early help with designing one of our robot domains, as well as helpful comments on an early draft of this paper. We acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the simulation results reported within this paper. Finally, we wish to thank our three Spot robots, Moana, Donner, and Kepler, for being so reliable throughout the extensive prototyping and experimentation required for this paper. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of our sponsors.